OpenStudy (liv33):

For a medal and fan, find the approximate side length of a square game board with an area of 160 in^2 to the nearest whole inch.

OpenStudy (mathstudent55):

What is the area of a square?

OpenStudy (liv33):

A = a^2

OpenStudy (mathstudent55):

Correct. If the side length is a, then the area is a^2. 

OpenStudy (mathstudent55):

We know the area, and we know the formula, so we can write: a^2 = A We are given that A = 160 in.^2, so we replace A with 160 in.^2 to get a^2 = 160 in.^2 Now we need to find a.

OpenStudy (liv33):

That makes sense

OpenStudy (mathstudent55):

Yes, believe it or not, math does make sense. We have the square of a, but we want a, so we take the square root of both sides. \(a^2 = 160~in.^2\) \(\sqrt{a^2} = \sqrt{160 ~in.^2}\) \(a^2 = \sqrt{160 ~in.^2}\) Now you need to take the square root of 160. Do you have a calculator with square root?

OpenStudy (liv33):

Square root of 160 is 12.650, rounded

OpenStudy (mathstudent55):

Since we need the side length rounded to the nearest whole number, we don't even need a calculator. Let's start with 10^2 which we know is 100. Then 11^2 = 121 too small 12^2 = 144 too small 13^2 = 169 too big The square root of 160 must be between 12 and 13, but closer to 13 than to 12.

OpenStudy (liv33):

Right